Predict the Normality of H2SO4 if 75 ml of 96.6 % pure H2SO4 added to 425 ml water. The density of H2SO4 is 1.83 g/cm3?
1
Expert's answer
2021-07-14T05:05:12-0400

M = n/V where n is the number of moles and V is the volume in litres,

d = density = 1.83 g/ml = 1830 g/L,

ie. weight of H2SO4 is 1830 g in 1 L of solution,

But H2SO4 is present 96.6% by weight in the solution i.e 0.966 x 1830 = 1767.78 g,

So M = n/V = 1767.78/98/1 where 98 is the molar mass of H2SO4,

= 18.039 m/L